OpenStudy (anonymous):

solve x+2y=2, 5x+3y=-18

OpenStudy (anonymous):

solve your first equation for x

OpenStudy (anonymous):

do you know how to do this @jasco29645 ?

OpenStudy (anonymous):

sure dont..i'm new at algebra

OpenStudy (anonymous):

ok well start by subtracting 2y from both sides of the first equation. what do you get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

you would get x=-2y+2 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so take that and substitute it for x in the second equation, what would that look like?

OpenStudy (anonymous):

5(0)+3y=-18

OpenStudy (anonymous):

where are you getting 0 from?

OpenStudy (anonymous):

you said to substitiute my previous answer for x for the 2nd equation

OpenStudy (anonymous):

i meant the equation that i gave you sorry

OpenStudy (anonymous):

0=-2y+2

OpenStudy (anonymous):

instead of 0 use x x=-2y+2

OpenStudy (anonymous):

now if x=-2y+2 then you can plug in 5(-2y+2)+3y=-18 instead of 5x+3y=-18

OpenStudy (anonymous):

you following me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know what the distributive property is?

OpenStudy (anonymous):

yes...so 3x/3=-18/3 so x=-6

OpenStudy (anonymous):

so now solve for y

OpenStudy (anonymous):

no the distributive property is x(a+b)=xa+xb

OpenStudy (anonymous):

5(-2y+2)+3y=-18 so with the distributive property we get -2y(5)+2(5)+3y=-18

OpenStudy (anonymous):

so what does that simplify too?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

because the -2y(5)=2(5) cancels out + 3==18 divided by 3 = - 6

OpenStudy (anonymous):

no you cant cancel out unlike terms you would get -10y+10+3y=-18

OpenStudy (anonymous):

so when you add like terms you get all the y's to one side and the other numbers to the other

OpenStudy (anonymous):

so -10y+3y=-7

OpenStudy (anonymous):

no you need to subtract 10 to both sides so you would get -10y+3y=-18-10

OpenStudy (anonymous):

so add -10y+3y, and add up -18-10

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

oops not negative..just 4

OpenStudy (anonymous):

-7y=-28 divide both sides by -7 and you get y=4

OpenStudy (anonymous):

you cant just forget about the y

OpenStudy (anonymous):

now that we have the y plug that back into the equation 5x+3y=-18

OpenStudy (anonymous):

so what would our first step be?

OpenStudy (anonymous):

@jasco29645 ?